# On the Robustness of Kolmogorov-Arnold Networks: An Adversarial Perspective

### v2

@@ Abstract
Kolmogorov-Arnold Networks (KANs) have recently emerged as a novel approach to function approximation, demonstrating remarkable potential in various domains. Despite their theoretical promise, the robustness of KANs under adversarial conditions has yet to be thoroughly examined. In this paper we explore the adversarial robustness of KANs, with a particular focus on image classification tasks. We assess the performance of KANs against standard white-box and black-box adversarial attacks, comparing their resilience to that of established neural network architectures. Our experimental evaluation encompasses a variety of standard image classification benchmark datasets and investigates both fully connected and convolutional neural network architectures, of three sizes: small, medium, and large. We conclude that small- and medium-sized KANs (either fully connected or convolutional) are not consistently more robust than their standard counterparts, but that large-sized KANs are, by and large, more robust. This comprehensive evaluation of KANs in adversarial scenarios offers the first in-depth analysis of KAN security, laying the groundwork for future research in this emerging field.
@@ 1 Introduction
This paper delves into the robustness of KANs when faced with adversarial attacks under white-box and black-box conditions. Using the MNIST, KMNIST, FashionMNIST, CIFAR-10, and SVHN datasets, our analysis not only sheds light on the robustness of KANs but also contributes to the broader understanding of neural network security in adversarial settings.
The next section provides a comprehensive overview of the architectures employed in this study, accompanied by a detailed discussion of their robustness. Section 3 outlines the experimental setup used to evaluate our proposed approach. Section 4 examines the robustness of the different architectures against adversarial attacks. Finally, we present concluding remarks in Section 5.
@@ 2.2 Robustness in Neural Networks
Such examples underscore the critical need for developing networks with enhanced robustness against adversarial attacks. Numerous other attack methodologies further highlight this necessity, revealing the extensive landscape of adversarial vulnerability in neural networks (Goodfellow et al., 2014; Lapid et al., 2024b; Szegedy et al., 2013; Tamam et al., 2023; Wei et al., 2022; Lapid et al., 2022; Chen et al., 2024; Carlini & Wagner, 2017; Lapid & Sipper, 2023b; Andriushchenko et al., 2020; Lapid & Sipper, 2023a).
@@ 2.3 Adversarial Attack Types
Adversarial attacks are techniques used to evaluate the robustness of neural networks by introducing small perturbations to input data, causing the model to misclassify the input. These attacks are broadly categorized into white-box attacks and black-box attacks, based on the adversary’s level of access to the model.
White-Box Attacks. In white-box attacks, the adversary has full knowledge of the target model, including its architecture, weights, and gradients. This access allows the adversary to compute perturbations directly by exploiting the model’s vulnerabilities. White-box attacks represent the strongest adversarial threat model, as they assume complete transparency of the system.
Black-Box Attacks. In black-box attacks, the adversary has no direct access to the model’s internal parameters or gradients. Instead, the adversary generates adversarial examples by querying the model or leveraging transferability across architectures. Black-box attacks are more practical in real-world scenarios, as they mimic situations where the adversary does not have access to proprietary models or data.
@@ 3 Experimental Setup
Do KANs offer enhanced security compared to classical neural network architectures, when confronted with an adversary?
To address this question we conduct an experimental evaluation of KANs. While it may appear unconventional to begin with experiments, direct assessment of the security of KANs is most effectively achieved through targeted attacks and subsequent analyses. We evaluate the robustness of FCKANs, CKANs, FCNNs and CNNs under white-box and black-box adversarial attacks. Both fully connected and convolutional architectures are tested across multiple datasets, and the results are analyzed to compare the robustness of the two architectures.
@@ 3.2 Adversarial Attacks
Furthermore, we evaluated the robustness of all the models against three different black-box attack methodologies: transfer-based attacks (Section B.2.1), simple black-box attack (SimBA, (Tu et al., 2019)), square attack (Andriushchenko et al., 2020), and natural evolutionary strategies attack (NES, (Ilyas et al., 2018)). White-box attacks were evaluated on all test sets, while black-box attacks were evaluated on a random sample of 1,000 images from the corresponding test set. Transferability metrics were calculated using 5,000 images from the test sets.
The hyperparameter configurations for these experiments are provided in Table 3. For the grayscale datasets (MNIST, FashionMNIST, and KMNIST), an $\epsilon$ value of $32/255$ was utilized, whereas for the multi-channel; datasets (CIFAR-10 and SVHN), an $\epsilon$ value of $8/255$ was employed. All experiments were conducted with a batch size of 16.
Table 3: Hyperparameters used in experiments with white-box and black-box attacks.
Attack Type Attack Parameters (Grayscale Datasets) Parameters (Multi-channel Datasets)
@@
A comprehensive description of each attack can be found in Appendix B.
@@ 4.1 Fully Connected Models
Table 4: White-box and black-box attacks on fully connected models evaluated across MNIST, FashionMNIST, and KMNIST datasets. Robust accuracy is given for each corresponding attack. The “Acc” column refers to the clean accuracy of the model. Boldface values indicate the most robust model in each experiment.
Dataset Model White-Box Black-Box Acc
FGSM PGD C&W MIM Square SimBA NES
MNIST $\text{FCKAN}_{\text{small}}$ 6.96 0.00 0.01 0.00 0.09 0.00 0.00 96.02
$\text{FCNN}_{\text{small}}$ 1.40 0.52 0.65 0.62 2.38 0.19 2.08 97.94
$\text{FCKAN}_{\text{medium}}$ 9.51 0.00 0.00 0.00 0.00 0.19 0.00 97.62
$\text{FCNN}_{\text{medium}}$ 10.80 0.77 0.87 1.05 3.27 18.35 19.94 98.31
$\text{FCKAN}_{\text{large}}$ 37.83 1.64 2.27 0.42 24.20 38.78 0.00 94.54
$\text{FCNN}_{\text{large}}$ 9.57 4.78 0.15 4.76 2.38 5.75 5.25 97.47
FashionMNIST $\text{FCKAN}_{\text{small}}$ 4.84 0.00 0.00 0.00 0.00 0.00 0.00 86.83
$\text{FCNN}_{\text{small}}$ 0.95 0.23 0.44 0.32 3.86 0.39 3.67 88.20
$\text{FCKAN}_{\text{medium}}$ 3.78 0.00 0.00 0.00 0.00 0.00 0.00 86.72
$\text{FCNN}_{\text{medium}}$ 5.23 0.12 0.11 0.18 1.58 11.30 11.80 89.14
$\text{FCKAN}_{\text{large}}$ 16.68 0.05 0.09 0.00 19.64 29.06 0.00 83.24
$\text{FCNN}_{\text{large}}$ 1.48 0.11 0.00 0.11 0.19 2.48 2.67 88.53
KMNIST $\text{FCKAN}_{\text{small}}$ 2.80 0.00 0.00 0.00 0.00 0.00 0.00 81.71
$\text{FCNN}_{\text{small}}$ 3.09 1.16 1.52 1.41 7.24 0.89 5.55 90.28
$\text{FCKAN}_{\text{medium}}$ 7.65 0.00 0.00 0.00 0.00 1.48 1.68 87.47
$\text{FCNN}_{\text{medium}}$ 11.74 1.01 1.11 1.35 6.54 8.82 17.85 91.15
$\text{FCKAN}_{\text{large}}$ 25.90 0.36 0.53 0.04 14.78 23.61 0.00 78.92
$\text{FCNN}_{\text{large}}$ 7.14 0.70 0.29 1.05 3.86 1.38 2.67 89.28
@@
Figure 1: Comparative loss dynamics of FCNNs and FCKANs under PGD, MIM, and C&W attacks on the MNIST, FashionMNIST, and KMNIST datasets (from top to bottom, respectively). Each line represents the mean loss across batches per iteration, with shaded areas indicating the standard deviation.
@@
When considering black-box attacks (Square, SimBA, and NES), the performance dynamics shift. Small FCNNs outperform FCKANs on all datasets, where small FCKANs failed to classify any image generated with those attacks. When considering the medium models, although FCKANs successfully classified a small number of images (especially on the KMNIST dataset), medium FCNNs still performed better than medium FCKANs across all datasets. However, for larger models, FCKANs demonstrate a marked improvement in robustness against SimBA and Square black-box attacks, significantly surpassing FCNNs. In contrast, similar to the small and medium sizes under NES attack, FCKANs failed to classify any adversarial image. The fact that FCKANs failed under the NES attack could result from the gradient approximation for these architectures being much more accurate than for the FCNN architecture.
@@
Table 5: Maximum transferability of our fully connected models, for FashionMNIST, KMNIST, and MNIST. The value in row $i$, column $j$ represents the maximum transferability between row-$i$ model and column-$j$ model across four attacks—FGSM, PGD, C&W, MIM—calculated using Equation 11. Unlike robustness tables, the values here indicate the percentage of images misclassified. The ‘Average’ row shows the average transferability for the attacking model. Boldfaced values denote the lowest transferability, where a lower value indicates better robustness.
@@
$\text{FCKAN}_{\text{small}}$ 100.00 57.32 38.97 59.98 56.23 48.76
$\text{FCKAN}_{\text{medium}}$ 40.55 100.00 28.11 39.89 39.99 35.81
$\text{FCKAN}_{\text{large}}$ 32.34 29.91 100.00 31.90 26.03 27.18
$\text{FCNN}_{\text{small}}$ 81.34 81.57 59.43 99.76 98.46 90.64
$\text{FCNN}_{\text{medium}}$ 64.82 68.91 45.16 90.19 99.86 85.33
$\text{FCNN}_{\text{large}}$ 57.13 59.31 39.45 75.18 82.21 100.00
Average ($\downarrow$) 62.69 66.16 51.85 66.15 67.13 64.61
@@
$\text{FCKAN}_{\text{small}}$ 100.00 29.58 21.11 21.60 13.77 19.71
$\text{FCKAN}_{\text{medium}}$ 32.87 100.00 17.47 16.19 10.94 17.43
$\text{FCKAN}_{\text{large}}$ 19.59 11.82 99.94 7.99 4.69 11.30
$\text{FCNN}_{\text{small}}$ 70.68 62.29 45.95 98.72 89.77 71.58
$\text{FCNN}_{\text{medium}}$ 51.04 42.42 30.76 83.49 98.88 59.40
$\text{FCNN}_{\text{large}}$ 45.75 35.72 31.89 50.22 43.74 99.66
Average ($\downarrow$) 53.32 46.97 41.18 46.36 43.46 46.51
@@
$\text{FCKAN}_{\text{small}}$ 100.00 28.10 22.94 55.10 30.55 32.32
$\text{FCKAN}_{\text{medium}}$ 35.33 100.00 17.71 38.83 22.67 26.74
$\text{FCKAN}_{\text{large}}$ 22.78 08.94 99.76 18.50 08.36 14.56
$\text{FCNN}_{\text{small}}$ 81.35 68.48 48.68 99.78 95.84 88.31
$\text{FCNN}_{\text{medium}}$ 64.17 54.40 33.83 95.48 99.62 84.39
$\text{FCNN}_{\text{large}}$ 58.89 44.78 31.77 78.42 72.29 99.92
Average ($\downarrow$) 60.41 50.78 42.44 64.35 54.88 57.70

### v3

@@ Abstract
Kolmogorov-Arnold Networks (KANs) have recently emerged as a novel paradigm for function approximation by leveraging univariate spline-based decompositions inspired by the Kolmogorov–Arnold theorem. Despite their theoretical appeal—particularly the potential for inducing smoother decision boundaries and lower effective Lipschitz constants—their adversarial robustness remains largely unexplored. In this work, we conduct the first comprehensive evaluation of KAN robustness in adversarial settings, focusing on both fully connected (FCKANs) and convolutional (CKANs) instantiations for image classification tasks. Across a wide range of benchmark datasets (MNIST, FashionMNIST, KMNIST, CIFAR-10, SVHN, and a subset of ImageNet), we compare KANs against conventional architectures using an extensive suite of attacks, including white-box methods (FGSM, PGD, C&W, MIM), black-box approaches (Square Attack, SimBA, NES), and ensemble attacks (AutoAttack). Our experiments reveal that while small- and medium-scale KANs are not consistently more robust than their standard counterparts, large-scale KANs exhibit markedly enhanced resilience against adversarial perturbations. An ablation study further demonstrates that critical hyperparameters—such as number of knots and spline order—significantly influence robustness. Moreover, adversarial training experiments confirm the inherent safety advantages of KAN-based architectures. Overall, our findings provide novel insights into the adversarial behavior of KANs and lay a rigorous foundation for future research on robust, interpretable network designs.
@@ 1 Introduction
This paper delves into the robustness of KANs when faced with adversarial attacks under white-box and black-box conditions. Using multiple datasets—MNIST, KMNIST, FashionMNIST, CIFAR-10, SVHN, and a subset of ImageNet—our analysis not only sheds light on the robustness of KANs but also contributes to the broader understanding of neural network safety in adversarial settings.
The next section provides a comprehensive overview of the architectures employed in this study, along with a detailed discussion of their robustness. In Section 3 we describe the experimental setup and evaluation protocols used to assess our proposed approach. Section 4 then presents the experimental results, reporting how the different architectures perform under a variety of adversarial attacks. Finally, we conclude our study and outline future research directions in Section 5.
@@ 2.2 Robustness in Neural Networks
Such examples underscore the critical need for developing networks with enhanced robustness against adversarial attacks. Numerous other attack methodologies further highlight this necessity, revealing the extensive landscape of adversarial vulnerability in neural networks (Goodfellow et al., 2014; Lapid et al., 2024b; Szegedy et al., 2013; Tamam et al., 2023; Wei et al., 2022; Lapid et al., 2022; Chen et al., 2024; Carlini & Wagner, 2017; Lapid & Sipper, 2023; Andriushchenko et al., 2020; Lapid et al., 2024c).
@@
2.3 Theoretical Motivation for KAN
Although the primary thrust of our work is empirical, there are several theoretical underpinnings suggesting why KANs may enjoy enhanced adversarial robustness compared to classical neural architectures. In particular, two concepts are worthy of discussion: (1) the functional decomposition principle from Kolmogorov–Arnold theory, and (2) connections to Lipschitz constants and their role in adversarial vulnerability.
2.3.1 Kolmogorov-Arnold Function Decomposition
KANs draw inspiration from the Kolmogorov Superposition Theorem (Kolmogorov, 1961; Arnold, 2009), which states that any continuous multivariate function can be decomposed into a finite sum of compositions of continuous univariate functions. In a neural-network instantiation, this translates to learning univariate spline-based transformations, followed by mixing via simpler aggregations. Concretely, one approximates a target function $f\colon\mathbb{R}^{n}\to\mathbb{R}^{m}$ by expressing it as a sum or composition of univariate mappings (often denoted $\phi_{i}$ and $\psi_{i}$).
Such a decomposition imposes inherent structure on the learned function. Instead of learning a fully unconstrained, high-dimensional parameter space, KANs learn univariate blocks (or “knots”) that are each individually simpler. This structure may yield smoother learned decision boundaries, potentially mitigating the local sensitivity that conventional fully connected or convolutional layers can exhibit (Cisse et al., 2017; Ross & Doshi-Velez, 2018).
2.3.2 Lipschitz Perspective and Adversarial Vulnerability
A prominent explanation for adversarial vulnerability involves the Lipschitz constant of the model’s decision function: networks with larger Lipschitz constants are typically more sensitive to small local perturbations (Cisse et al., 2017; Goodfellow et al., 2014). Formally, a function $f$ is $L$-Lipschitz if:
$\|f(x)-f(x^{\prime})\|\leq L\cdot\|x-x^{\prime}\|\quad\text{for all }x,x^{\prime}\in\mathbb{R}^{n}.$ (1)
When $L$ is large, small input perturbations can lead to disproportionally large changes in the output (Goodfellow et al., 2014), facilitating adversarial attacks.
In KANs, the piecewise or univariate spline decomposition can, in principle, distribute the function’s complexity across multiple univariate components. This might help control or lower effective Lipschitz constants, as each univariate block need not learn steep or extreme gradients. Instead, the overall function is composed of smaller, potentially smoother parts. Empirically, this could translate into fewer local minimas where tiny perturbations trigger large output changes.
@@ 2.4 Adversarial Attack Types
Adversarial attacks are techniques used to evaluate the robustness of neural networks by introducing small perturbations to input data, causing the model to misclassify the input (Szegedy et al., 2013; Goodfellow et al., 2014). These attacks are broadly categorized into white-box attacks and black-box attacks, based on the adversary’s level of access to the model.
White-Box Attacks. In white-box attacks, the adversary has full knowledge of the target model, including its architecture, weights, and gradients (Goodfellow et al., 2014; Madry et al., 2017). This access allows the adversary to compute perturbations directly by exploiting the model’s vulnerabilities. White-box attacks represent the strongest adversarial threat model, as they assume complete transparency of the system.
Black-Box Attacks. In black-box attacks, the adversary has no direct access to the model’s internal parameters or gradients (Ilyas et al., 2018; Tu et al., 2019). Instead, the adversary generates adversarial examples by querying the model or leveraging transferability across architectures (Szegedy et al., 2013). Black-box attacks are more practical in real-world scenarios, as they mimic situations where the adversary does not have access to proprietary models or data.
Ensemble-Based Attacks. In addition to individual white-box and black-box attacks, ensemble-based attacks combine multiple adversarial strategies to evaluate model robustness more reliably (Croce & Hein, 2020).
@@ 3 Experimental Setup
Do KANs offer enhanced safety compared to classical neural network architectures, when confronted with an adversary?
To address this question we conduct an experimental evaluation of KANs. While it may appear unconventional to begin with experiments, direct assessment of the safety of KANs is most effectively achieved through targeted attacks and subsequent analyses. We evaluate the robustness of FCKANs, CKANs, FCNNs and CNNs under white-box and black-box adversarial attacks. Both fully connected and convolutional architectures are tested across multiple datasets, and the results are analyzed to compare the robustness of the two architectures.
@@ 3.1 Classifier Models
In addition to the standard datasets, we trained convolutional architecutres on a 10-image subset of ImageNet (Deng et al., 2009) to evaluate their performance on high-resolution images. Due to memory constraints, the input resolutions varied based on model size: Small CKAN & CNN models trained on 224 × 224 images, Medium CKAN & CNN models trained on 160 × 160 images, and Large CKAN & CNN models trained on 112 × 112 images. The learning rate, optimizer, and weight decay were kept consistent with other convolutional models, but the batch size was reduced to 4.
@@ 3.2 Adversarial Attacks
Furthermore, we evaluated the robustness of all the models against three different black-box attack methodologies: transfer-based attacks (Section B.2.1), simple black-box attack (SimBA, (Tu et al., 2019)), square attack (Andriushchenko et al., 2020), and natural evolutionary strategies attack (NES, (Ilyas et al., 2018)). We also added AutoAttack (Croce & Hein, 2020), an ansemble of adversarial attacks including PGD, FAB, Square Attack, and a verification-based attack, which provides a comprehensive robustness evaluation. White-box attacks and AutoAttack were evaluated on all test sets, while black-box attacks were evaluated on a random sample of 1,000 images from the corresponding test set. Transferability metrics were calculated using 5,000 images from the test sets.
The hyperparameter configurations for these experiments are provided in Table 4. For the grayscale datasets (MNIST, FashionMNIST, and KMNIST), an $\epsilon$ value of $32/255$ was utilized, whereas for the multi-channel datasets (CIFAR-10, SVHN, and ImageNet), an $\epsilon$ value of $8/255$ was employed. All experiments were conducted with a batch size of 16, except for ImageNet, which was trained with a batch size of 4. Furthermore, in our implementation of AutoAttack, we adhered to the default hyperparameter settings, as specified in the original AutoAttack paper (Croce & Hein, 2020).
Table 3: Hyperparameters used in experiments with white-box attacks, black-box attacks and ensemble-based attacks. $k$: number of iterations; $\alpha$: step size (learning rate) for perturbations; $c$: confidence parameter in the C&W attack; $\mu$: momentum factor in MIM; $\sigma$: standard deviation of the noise in NES; $n$: number of samples used in NES to estimate gradients. For more details see Appendix B.
Attack Type Attack Params (Grayscale Datasets) Params (Multi-channel Datasets)
@@
Table 4: Ablation study configurations: Varying number of knots and spline order.
Config # Knots Spline Order
A 3 5
B 3 7
C 5 5
D 5 3
E 7 3
3.3 Ablation Study: Impact of KAN Hyperparameters on Robustness
To investigate how KAN hyperparameters affect adversarial robustness, we performed an ablation study focusing on two key factors: number of knots and spline order. These parameters directly influence the expressiveness and smoothness of function approximation in KANs, which may affect their susceptibility to adversarial attacks. We systematically varied these hyperparameters across all sizes of FCKAN models to assess their impact on adversarial robustness. Each model was trained and evaluated using the same attack settings as described in Section 3.2. The five configurations tested are delineated in Table 4.
3.4 Adversarial training
To enhance model robustness, we conducted adversarial training using PGD attack. Adversarial training is a defense strategy where models are trained on adversarially perturbed examples to improve their resilience against future attacks. We applied adversarial training to all the sizes of both fully connected and conv architectures, selecting different datasets for each. The FCKANs and FCNNs were trained on KMNIST data set. The CKANs and CNNs were trained on the CIFAR-10 dataset. Each model was trained for 20 epochs, using a batch size of 16. To asses the effectiveness of adversarial training, we evaluated the trained models against all the white-box and black-box attacks listed in Table 4.
@@ 4.1 Fully Connected Models
Table 5: White-box attacks, black-box attacks, and ensemble attacks on fully connected models, evaluated across MNIST, FashionMNIST, and KMNIST datasets. Robust accuracy is given for each corresponding attack. The “Acc” column refers to the clean accuracy of the model. Boldface values indicate the most robust model in each experiment.
Dataset Model White-Box Black-Box Ensemble-Based Acc
FGSM PGD C&W MIM Square SimBA NES Auto Attack
MNIST $\text{FCKAN}_{\text{small}}$ 6.96 0.00 0.01 0.00 0.09 0.00 0.00 0.00 96.02
$\text{FCNN}_{\text{small}}$ 1.40 0.52 0.65 0.62 2.38 0.19 2.08 0.30 97.94
$\text{FCKAN}_{\text{medium}}$ 9.51 0.00 0.00 0.00 0.00 0.19 0.00 0.00 97.62
$\text{FCNN}_{\text{medium}}$ 10.80 0.77 0.87 1.05 3.27 18.35 19.94 0.25 98.31
$\text{FCKAN}_{\text{large}}$ 37.83 1.64 2.27 0.42 24.20 38.78 0.00 0.00 94.54
$\text{FCNN}_{\text{large}}$ 9.57 4.78 0.15 4.76 2.38 5.75 5.25 0.05 97.47
FashionMNIST $\text{FCKAN}_{\text{small}}$ 4.84 0.00 0.00 0.00 0.00 0.00 0.00 0.00 86.83
$\text{FCNN}_{\text{small}}$ 0.95 0.23 0.44 0.32 3.86 0.39 3.67 0.14 88.20
$\text{FCKAN}_{\text{medium}}$ 3.78 0.00 0.00 0.00 0.00 0.00 0.00 0.00 86.72
$\text{FCNN}_{\text{medium}}$ 5.23 0.12 0.11 0.18 1.58 11.30 11.80 0.01 89.14
$\text{FCKAN}_{\text{large}}$ 16.68 0.05 0.09 0.00 19.64 29.06 0.00 0.00 83.24
$\text{FCNN}_{\text{large}}$ 1.48 0.11 0.00 0.11 0.19 2.48 2.67 0.00 88.53
KMNIST $\text{FCKAN}_{\text{small}}$ 2.80 0.00 0.00 0.00 0.00 0.00 0.00 0.00 81.71
$\text{FCNN}_{\text{small}}$ 3.09 1.16 1.52 1.41 7.24 0.89 5.55 0.90 90.28
$\text{FCKAN}_{\text{medium}}$ 7.65 0.00 0.00 0.00 0.00 1.48 1.68 0.00 87.47
$\text{FCNN}_{\text{medium}}$ 11.74 1.01 1.11 1.35 6.54 8.82 17.85 0.37 91.15
$\text{FCKAN}_{\text{large}}$ 25.90 0.36 0.53 0.04 14.78 23.61 0.00 0.00 78.92
$\text{FCNN}_{\text{large}}$ 7.14 0.70 0.29 1.05 3.86 1.38 2.67 0.06 89.28
@@
Figure 1: Comparative loss dynamics of FCNNs and FCKANs under PGD, MIM, and C&W attacks on the MNIST, FashionMNIST, and KMNIST datasets (from top to bottom, respectively). Each line represents the mean loss across batches, per iteration, with shaded areas indicating the standard deviation.
@@
When considering black-box attacks (Square, SimBA, and NES), the performance dynamics shift. Small FCNNs outperform FCKANs on all datasets, where small FCKANs failed to classify any image generated with those attacks. When considering the medium models, although FCKANs successfully classified a small number of images (especially on the KMNIST dataset), medium FCNNs still performed better than medium FCKANs across all datasets. However, for larger models, FCKANs demonstrate marked improvement in robustness against SimBA and Square black-box attacks, significantly surpassing FCNNs. In contrast, similar to the small and medium sizes under the NES attack, FCKANs failed to classify any adversarial image. The fact that FCKANs failed under the NES attack could result from the gradient approximation for these architectures being much more accurate than for the FCNN architecture.
All the FCNN models, except the large FCNN on FashionMNIST, successfully classify at least one image generated with AutoAttack, while none of the FCKAN models, across all datasets, successfully classify an image generated with that attack. This may be due to the structural properties of FCKANs, which, while beneficial for robustness against certain attacks, might also introduce vulnerabilities that AutoAttack effectively exploits. Specifically, AutoAttack’s adaptive nature could be particularly well-suited to finding perturbations that significantly degrade FCKAN performance.
@@
Table 6: Maximum transferability of our fully connected models, for FashionMNIST, KMNIST, and MNIST. Each row represents the model that generates the adversarial example and each column represents the model that evaluates those examples. The value in row $i$, column $j$ represents the maximum transferability between row-$i$ model and column-$j$ model across four attacks—FGSM, PGD, C&W, MIM—calculated using Equation 12. Unlike robustness tables, the values here indicate the percentage of images misclassified. The ‘Average’ row shows the average transferability for the attacking model. Boldfaced values denote the lowest transferability, where a lower value indicates better robustness.
@@
$\text{FCKAN}_{\text{small}}$ 57.32 38.97 59.98 56.23 48.76
$\text{FCKAN}_{\text{medium}}$ 40.55 - 28.11 39.89 39.99 35.81
$\text{FCKAN}_{\text{large}}$ 32.34 29.91 - 31.90 26.03 27.18
$\text{FCNN}_{\text{small}}$ 81.34 81.57 59.43 - 98.46 90.64
$\text{FCNN}_{\text{medium}}$ 64.82 68.91 45.16 90.19 - 85.33
$\text{FCNN}_{\text{large}}$ 57.13 59.31 39.45 75.18 82.21 -
Average ($\downarrow$) 55.23 59.40 42.22 59.42 60.58 57.54
@@
$\text{FCKAN}_{\text{small}}$ 29.58 21.11 21.60 13.77 19.71
$\text{FCKAN}_{\text{medium}}$ 32.87 - 17.47 16.19 10.94 17.43
$\text{FCKAN}_{\text{large}}$ 19.59 11.82 - 7.99 4.69 11.30
$\text{FCNN}_{\text{small}}$ 70.68 62.29 45.95 - 89.77 71.58
$\text{FCNN}_{\text{medium}}$ 51.04 42.42 30.76 83.49 - 59.40
$\text{FCNN}_{\text{large}}$ 45.75 35.72 31.89 50.22 43.74 -
Average ($\downarrow$) 43.98 36.36 29.43 35.89 32.58 35.88
@@
$\text{FCKAN}_{\text{small}}$ 28.10 22.94 55.10 30.55 32.32
$\text{FCKAN}_{\text{medium}}$ 35.33 - 17.71 38.83 22.67 26.74
$\text{FCKAN}_{\text{large}}$ 22.78 08.94 - 18.50 08.36 14.56
$\text{FCNN}_{\text{small}}$ 81.35 68.48 48.68 - 95.84 88.31
$\text{FCNN}_{\text{medium}}$ 64.17 54.40 33.83 95.48 - 84.39
$\text{FCNN}_{\text{large}}$ 58.89 44.78 31.77 78.42 72.29 -
Average ($\downarrow$) 52.50 40.94 30.98 57.26 45.94 49.26